# MrGuard: A Multilingual Reasoning Guardrail for Universal LLM Safety

### v2

@@ Abstract
Dan Roth University of Pennsylvania danr@seas.upenn.edu                        Insup Lee University of Pennsylvania lee@seas.upenn.edu
@@ 1 Introduction
To address these issues, safety alignment techniques—such as Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO)—have been proposed to align LLM behavior with human values and reduce unsafe generations (Ouyang et al., 2022; Rafailov et al., 2023). Another line of work focuses on building standalone safety classifiers or guardrails, which act as filters to detect and block unsafe user prompts or model generations without modifying the LLM itself (Inan et al., 2023; Ghosh et al., 2024). These lightweight safety modules are advantageous in being more efficient and easier to deploy or update (Table 1). Most existing methods are primarily English-centric (Liu et al., 2025; Kang and Li, 2024b; Yuan et al., 2024) which cannot handle multilingual content moderation (Yang et al., 2024). As shown in Figure 1, the guardrail model successfully identifies the unsafe user prompt in English but fails to detect its semantically equivalent counterpart in Chinese. Moreover, without explanations, it becomes difficult to understand the rationale behind the guardrail’s decisions [^1] We interchangeably use guard and guardrail through the paper..
@@
To bridge this gap, our work is the first one to focus on building a guardrail tailored for multilingual safety scenarios with reasoning ability. We aim to design a robust, reasoning-aware safety guardrail that can effectively moderate harmful prompts across diverse languages and cultural contexts. Our contributions can be listed as follows:
@@
  • We achieve state-of-the-art results on several multilingual safety benchmarks, outperforming all baselines in prompt classification accuracy. We further demonstrate MR.Guard’s robustness on multilingual variations such as code-switching and sandwich Upadhayay and Behzadan (2024) attacks. Our results show that post-training with reasoning abilities significantly improves the robustness and performance of guardrails on multilingual prompt classification.
  • We present a comprehensive evaluation of our reasoning-enhanced guardrail with key metrics such as cross-lingual consistency and reasoning fidelity which establishes a strong baseline for future assessments of guardrail reasoning capabilities.
@@ 3.3 Curriculum-based GRPO
This reward penalizes formatting errors. If the output does not contain a properly formatted safety prediction (i.e., "Safety: safe" or "Safety: unsafe") which often happens in multilingual generation, the reward is $-1$. Otherwise, the reward is $1$.
@@
If the safety prediction is correct, the reward is $1$, otherwise, the reward is $-1$.
@@ 4.1 Experimental Setup
In our experiments, we use the training set from Aegis-2.0-Safety (Ghosh et al., 2025) as the English seed data. Our base model is LLaMA-3.1-8B-Instruct (Aaron Grattafiori, 2024), and we apply QLoRA (Dettmers et al., 2023) for parameter-efficient fine-tuning during both the SFT and GRPO stage [^3] More details on the training hyperparameters and configurations are in Appendix B.. To construct the curriculum using back-translation, we employ the facebook/nllb-200-3.3B model for translation and use all-MiniLM-L6-v2 to compute the sentence embeddings of both the original and the back-translated sentences. For the difficulty threshold, we set $t_{1}=0.85$ and $t_{2}=0.7$. To determine the language of the sampled output, we utilize an xlm-based language detector  [^4] The full name of the language detector: papluca/xlm-roberta-base-language-detection. Our experiments divide the test sets into two categories: in-domain languages, which are covered during training, and out-of-domain languages, which are not seen during training.
@@
MrGuard (Ours) 90.95 86.90 98.18 98.11 94.56 92.06 91.45 92.68 97.41 96.15
Table 2: Performance of different guardrails to identify multilingual safety across five benchmark datasets. We report F1 scores as the evaluation metric and bold the best-performing results for each dataset where ID refers to in-domain languages and OOD refers to out-of-domain languages. Top are baselines and the bottom part is MrGuard (ours). The model size and training dataset size are listed in Table 1.
@@ 4.2 Multilingual Content Moderation
Moreover, we present a language-wise breakdown in Figure 3 with more results in Appendix B. Our model consistently outperforms the baselines across a wide range of languages, although some baselines achieve comparable results on the English subset. We also observe that training on English-only data negatively impacts generalization to non-English languages. For example, although Aegis-2.0 and WildGuard perform well on English inputs, their performance degrades significantly on non-English data, whereas MrGuard maintains high performance and shows minimal performance drop, even on unseen languages. This further demonstrates the robustness and cross-lingual generalization capabilities of our model.
@@ 6 Conclusion
Potential Bias We use a single LLM as a judge to verify safety labels of translations, which may introduce bias inherent to the LLM. Additionally, our current evaluation of reasoning coherence and faithfulness between explanations and final safety predictions relies on automated heuristics, which may not perfectly align with human judgments.
@@ Appendix B Experiment Setup
We use the TRL library (von Werra et al., 2020) for both SFT and GRPO stage. For both stages, we set the LoRA rank and alpha to 32, with a dropout rate of 0.1. During SFT, we use a learning rate of SFT is $52e-$ 5 and train for 3 epochs. For GRPO, we set the learning rate to $51e-$ 5 and the number of training epoch is 1. We conduct a hyperparameter sweep over LoRA rank and alpha values $\{8,16,32\}$, and select the best configuration based on performance on the Aegis-2.0 validation dataset. The results are the average of three training runs. All use of the packages and artifacts are consistent with their intended use and license. We used ChatGPT to refine short sentences and paragraphs and to check for grammar errors.
@@ Appendix C Dataset Details
RTP-LX Aya MultiJail PTP_wildchat XSafety
ID EN, AR, ES, RU, ZH-Hans EN, AR, ES, RU EN, AR, ES, ZH EN, AR, ES, RU, ZH EN, AR, ES, RU, ZH
@@
Table 7: In-domain (ID) and out-of-domain (OOD) language coverage for each evaluation dataset.
@@ C.2 Human evaluation
We validated our LLM-generated synthetic translations (Section 3) by comparing the model’s safety judgments with human judgments. We have subsampled 100 sentences for ZH, RU, and AR. For each language, we have one volunteer, and the volunteers are native speakers in the target languages and are proficient in English. The guidelines of human evaluation is shown in Figure 8, and the results are listed in Table 9. The safety preservation rate is a three-point scale measurement (0-2). 0: The safety label is not preserved; 1: the safety label is preserved, but meaning is altered; 2: both safety label and original meaning are preserved.
@@ Appendix G Additional Results
We additionally perform code-switching and sandwich attack on XSafety dataset and the results are shown in Table 11 and Table 12.
Models EN $\uparrow$ Avg-CSRT $\uparrow$ $\Delta$downarrow
DUO-Guard 75.71 70.19 5.52
GuardR 85.99 84.42 1.57
LlaMa-Guard-3 64.48 61.08 3.40
Aegis-2.0 66.22 39.30 26.92
Wildguard 91.40 84.40 7.00
MrGuard 92.71 91.51 1.21
@@
G.1 Qualitative Result

### v3

@@ Abstract
Dan Roth University of Pennsylvania Oracle AI danr@seas.upenn.edu          Insup Lee University of Pennsylvania lee@seas.upenn.edu
@@ 1 Introduction
To address these challenges, safety alignment strategies, most notably Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO), which aim to align the behavior of LLMs with human values and thereby mitigate the risk of unsafe or harmful outputs (Ouyang et al., 2022; Rafailov et al., 2023). Another line of work focuses on building standalone safety classifiers or guardrails, which act as filters to detect and block unsafe user prompts or model generations without modifying the LLM itself (Inan et al., 2023; Ghosh et al., 2024). These lightweight safety modules are advantageous in being more efficient and easier to deploy or update (Table 1). Most existing methods are primarily English-centric (Liu et al., 2025; Kang and Li, 2024b; Yuan et al., 2024) which cannot handle multilingual content moderation (Yang et al., 2024). As shown in Figure 1, the guardrail model successfully identifies the unsafe user prompt in English but fails to detect its semantically equivalent counterpart in Chinese. Moreover, without explanations, it becomes difficult to understand the rationale behind the guardrail’s decisions [^1] We interchangeably use guard and guardrail through the paper..
@@
To bridge this gap, our work is the first one to focus on building a guardrail tailored for multilingual safety scenarios with reasoning ability. We aim to design a robust, reasoning-aware safety guardrail that can effectively moderate harmful prompts across diverse languages and cultural contexts. Our contributions can be listed as follows  [^2] Our code is available at https://github.com/yangy96/mrguard:
@@
  • We achieve state-of-the-art results on several multilingual safety benchmarks, outperforming all baselines in prompt classification accuracy. We further demonstrate MrGuard’s robustness on multilingual variations such as code-switching and sandwich Upadhayay and Behzadan (2024) attacks. Our results show that post-training with reasoning abilities significantly improves the robustness and performance of guardrails on multilingual prompt classification.
  • We present a comprehensive evaluation of our reasoning‐enhanced guardrail with key metrics such as cross‐lingual consistency and reasoning fidelity which establishes a strong baseline for future assessments of guardrail reasoning capabilities.
@@ 3.3 Curriculum-based GRPO
This reward penalizes formatting errors. If the output does not contain a properly formatted safety prediction (i.e., "Safety: safe" or "Safety: unsafe") which often happens in multilingual generation, the reward is 1-1. Otherwise, the reward is $1$.
@@
If the safety prediction is correct, the reward is $1$, otherwise, the reward is 1-1.
@@ 4.1 Experimental Setup
In our experiments, we use the training set from Aegis-2.0-Safety (Ghosh et al., 2025) as the English seed data. Our base model is LLaMA-3.1-8B-Instruct and LLaMA-3.2-3B-Instruct (Aaron Grattafiori, 2024), and we apply QLoRA (Dettmers et al., 2023) for parameter-efficient fine-tuning during both the SFT and GRPO stage [^4] More details on the training hyperparameters and configurations are in Appendix B.. To construct the curriculum using back-translation, we employ the facebook/nllb-200-3.3B model for translation and use all-MiniLM-L6-v2 to compute the sentence embeddings of both the original and the back-translated sentences. For the difficulty threshold, we set $t_{1}=0.85$ and $t_{2}=0.7$. To determine the language of the sampled output, we utilize an xlm-based language detector  [^5] The full name of the language detector: papluca/xlm-roberta-base-language-detection. Our experiments divide the test sets into two categories: in-domain languages, which are covered during training, and out-of-domain languages, which are not seen during training.
@@
MrGuard (8B) 91.04 86.32 98.16 98.21 94.33 92.06 91.17 92.15 97.26 95.74
MrGuard (3B) 88.04 85.13 95.44 94.73 91.89 90.60 88.22 88.56 95.09 89.40
Table 2: Performance of different guardrails to identify multilingual safety across five benchmark datasets. We report F1 scores as the evaluation metric and bold the best-performing results for each dataset where ID refers to in-domain languages and OOD refers to out-of-domain languages. Top are baselines and the bottom part is MrGuard (Ours) 8B and 3B. The model size and training dataset size are listed in Table 1.
@@ 4.2 Multilingual Content Moderation
Moreover, we present a language-wise breakdown in Figure 3 with more results in Appendix B [^7] We also include the breakdown of performance by the language family, language scripts and resource availability in the Appendix G.1.. Our model consistently outperforms the baselines across a wide range of languages, although some baselines achieve comparable results on the English subset. We also observe that training on English-only data negatively impacts generalization to non-English languages. For example, although Aegis-2.0 and WildGuard perform well on English inputs, their performance degrades significantly on non-English data, whereas MrGuard maintains high performance and shows minimal performance drop, even on unseen languages. This further demonstrates the robustness and cross-lingual generalization capabilities of our model.
@@ 5.3 Quality of Reasoning
We additionally perform human evaluation of the quality of the MrGuard’s generated reasoning. From the RTP_LX dataset, we subsampled 100 examples and generated reasoning in English, Chinese, and Hindi. Human volunteers then evaluated fidelity using the same instructions provided to GPT-4o in Figure 17. We show that the reasoning remain high fidelity on these languages in Table 7 (Here Hindi is an unseen and relatively-low resource language). We also conducted human evaluation (Appendix C.2) on safety preservation of syntactic data generation.
% EN ZH HI
Fidelity 98.0 87.0 86.0
Table 7: Explanation fidelity rates for a sampled subset of the RTP_LX dataset, as given by human annotators.
Moreover, we perform a deeper analysis on the failure cases of MrGuard’s reasoning and predictions. Below is the refined taxonomy for categorizing reasoning failures: A. Ambiguity: Covers both lexical ambiguity and unclear referents (e.g., pronouns without clear antecedents); B. Cultural Uncertainty: Reliance on culture-specific knowledge or assumptions that may not hold universally; C. Unsupported Inference: Jumps to conclusions without necessary premises or overgeneralizes from a single instance; D. Logical Inconsistency: Contradicts earlier steps or violates basic inference rules; E. Misclassification: Flags risks not present in the prompt. We apply GPT-4.1-mini as an automated judge to analyze the reasoning chains in the RTP_LX dataset. We report the percentage breakdown of reasoning assigned to each category in Table 8. We observe that the model rarely generates logic inconsistent reasoning but sometimes flag risks not present in the inputs. We also provide some qualitative examples of each category in Figure 16.
% EN ZH AR JA HI
A 5.04 6.67 7.88 8.99 6.72
B 7.91 26.67 23.65 48.31 44.78
C 2.88 6.67 4.93 3.93 2.99
D 0.0 0.0 0.0 0.0 0.0
E 94.96 95.78 81.77 73.60 70.15
Table 8: The percentage breakdown of reasoning for incorrect safety prediction from MrGuard under our failure taxonomy. Note here HI and JA are the unseen languages at the training.
@@ 6 Conclusion
Potential Bias We use a single LLM (GPT-4o-mini) as a judge to verify safety labels of translations, which may introduce bias inherent to the LLM. We acknowledge that relying on a single LLM for both generation and evaluation raises reliability concerns. As future work, we will explore ensembles of multiple LLMs for multilingual synthetic data generation and evaluation. Additionally, our current evaluation of reasoning coherence and faithfulness between explanations and final safety predictions relies on automated heuristics, which may not perfectly align with human judgments. Human Annotation To validate our synthetic data and the fidelity of LLM-generated reasoning, we conducted a small human-evaluation study on a subsampled dataset. Due to verification costs, we could not scale to multiple annotators or a larger sample size.
@@
Acknowledgment
We thank the anonymous reviewers for their constructive feedback and insightful suggestions. We would also like to thank Dr. Oleg Sokolsky and Dr. Almiqdad Saeed for their help with the synthetic data evaluation. Research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-20-1-0080. The views expressed are those of the authors and do not reflect the official policy or position of the Army Research Office or the U.S. Government.
@@ Appendix B Experiment Setup
We use the TRL library (von Werra et al., 2020) for both SFT and GRPO stage. For both stages, we set the LoRA rank and alpha to 32, with a dropout rate of 0.1. During SFT, we use a learning rate of SFT is 2e52e-5 and train for 3 epochs. For GRPO, we set the learning rate to 1e51e-5 and the number of training epoch is 1. We conduct a hyperparameter sweep over LoRA rank and alpha values $\{8,16,32\}$, and select the best configuration based on performance on the Aegis-2.0 validation dataset. All use of the packages and artifacts are consistent with their intended use and license. We used ChatGPT to refine short sentences and paragraphs and to check for grammar errors.
@@ Appendix C Dataset Details
RTP-LX Aya MultiJail Wildchat XSafety
ID EN, AR, ES, RU, ZH EN, AR, ES, RU EN, AR, ZH EN, AR, ES, RU, ZH EN, AR, ES, RU, ZH
@@
Table 9: In-domain (ID) and out-of-domain (OOD) language coverage for each evaluation dataset. For RTP_LX dataset, we used simplified Chinese version (ZH-Hans).
@@ C.2 Human evaluation
We validated our LLM-generated synthetic translations (Section 3) by comparing the model’s safety judgments with human judgments. We have subsampled 100 sentences for ZH, RU, and AR. For each language, we have one volunteer for each language, and the volunteers are native speakers in the target languages and are proficient in English. The guidelines of human evaluation is shown in Figure 8, and the results are listed in Table 11. The safety preservation rate is a three-point scale measurement (0-2). 0: The safety label is not preserved; 1: the safety label is preserved, but meaning is altered; 2: both safety label and original meaning are preserved.
@@ Appendix G Additional Results
G.1 Granular Breakdown of Performance
Here is the granular breakdown of performance by the language script/family/resource availability across different datasets.
By language script (Table 13)
Latin script: French, Spanish, Swahili, Italian, English, Serbian
Cyrillic script: Russian
Devanagari script: Hindi
Arabic script: Arabic
Hangul: Korean
Han script: Chinese
Japanese scripts: Japanese
Models Latin Cyrillic Devanagari Arabic Hangul Han Japanese
DUO-Guard 62.56 29.99 50.88 35.40 13.06 73.57 58.63
Guardreasoner 79.48 83.78 78.08 72.48 90.43 84.24 80.09
LlaMa-guard-3 67.16 64.92 59.54 62.73 77.43 62.25 66.24
Aegis-2.0 41.85 48.88 48.79 19.05 20.00 46.74 39.22
Wildguard 64.35 73.34 44.75 29.52 70.90 77.61 69.78
Ours 93.38 95.12 90.53 92.71 97.23 93.24 92.26
Table 13: Performance of different guardrails to identify multilingual safety across five benchmark datasets grouped by language script.
By language family (Table 14)
Afro-Asiatic: Arabic
Indo-European: French, Spanish, Italian, English, Russian, Serbian, Hindi
Sino-Tibetan: Chinese
Japonic: Japanese
Koreanic: Korean
Niger-Congo: Swahili
Models Indo-European Afro-Asiatic Sino-Tibetan Japonic Koreanic Niger-Congo
DUO-Guard 60.39 35.40 73.57 58.63 13.06 33.46
GuardR 86.69 72.48 84.24 80.09 90.43 31.90
LlaMa-guard-3 69.09 62.73 62.25 66.24 77.43 43.75
Aegis-2.0 49.20 19.05 46.74 39.22 20.00 4.37
Wildguard 71.30 29.52 77.61 69.78 70.90 5.08
Ours 95.20 92.71 93.24 92.26 97.23 79.48
Table 14: Performance of different guardrails to identify multilingual safety across five benchmark datasets grouped by language family.
By resource availability (Table 15)
We group languages according to their web‐coverage percentages as reported by Common Crawl.
High-Resource (>= 1%): English, Russian, Chinese, Spanish, French, Italian, Japanese
Mid-Resource (0.5% - 1%): Korean, Arabic
Low-Resource (<= 0.5%): Hindi, Serbian, Swahili.
Models High Mid Low
DUO-Guard 69.54 24.23 33.87
Guardreasoner 87.03 81.45 64.63
LlaMa-guard-3 68.25 70.08 59.36
Aegis-2.0 51.02 19.52 25.86
Wildguard 79.51 50.21 31.67
Ours 94.77 94.97 89.33
Table 15: Performance of different guardrails to identify multilingual safety across five benchmark datasets grouped by resource availability.
@@ G.1 Granular Breakdown of Performance
G.2 Additional Unseen Languages
Here we report results in Figure 14 on unseen mid/low-resource languages for different datasets. RTP_LX: Hebrew (HE); Aya: Filipino (FIL); XSafety: Bengali (BN); PTP_Wildchat: Korean (KO); MultiJail: Bengali (BN). We observe that our multilingual guardrail consistently outperforms the baselines.
Figure 14: F1 score breakdown on additional unseen languages from different datasets.
G.3 Additional Results on Multilingual Perturbations
We also perform code-switching and sandwich attack on XSafety dataset and the results are shown in Table 16 and Table 17.
Models EN $\uparrow$ Avg-CSRT $\uparrow$ $\Delta$ $\downarrow$
DUO-Guard 75.71 67.56 8.15
GuardR 85.99 84.41 1.58
LlaMa-Guard-3 64.48 60.81 3.67
Aegis-2.0 66.22 40.93 25.29
Wildguard 91.40 81.76 5.99
MrGuard 93.00 92.44 0.56
@@ G.3 Additional Results on Multilingual Perturbations
G.3.1 Additional Results on Hyperparameter Search
Here we provide additional results of model trained with different difficulty thresholds in Table 18.
($t_{2},t_{1}$) RTP_LX Aya Xsafety PTP MultiJ
(0.6,0.7) 90.46 97.56 92.73 90.16 96.49
(0.6,0.8) 90.69 97.67 93.32 90.68 97.38
(0.6,0.9) 90.61 97.64 92.66 90.48 96.69
(0.7,0.8) 90.81 97.75 93.18 90.82 96.85
(0.7,0.9) 90.16 97.20 92.81 90.33 96.74
(0.8,0.9) 90.13 97.30 91.91 89.83 96.57
Table 18: F1 scores on in-domain languages across datasets for models trained with varying difficulty thresholds.
G.4 Additional Results on Ablation Study
@@
We here reported the language match rate of the model trained without $R_{lang}$ to show the motivation of including the language component in the loss function in Table 20. As we show in the Table 5, our proposed model has a comparable performance with a model trained without $R_{lang}$ but the generated reasoning is not in the same language as the input.
EN AR ZH RU HI
Ours 97.30 98.76 99.52 98.86 99.91
W/o Rlang 97.40 0.0 0.0 0.0 6.64
Table 20: Language match rate of the model trained with/without $R_{lang}$ in the loss function.
Appendix H Qualitative Results
@@
We have defined a taxonomy of failure cases and provide qualitative examples of each case from different languages.
Figure 16: We present qualitative examples across languages illustrating our failure taxonomy. Inputs are from RTP_LX; analysis are generated by MrGuard. English translations via Google Translate are provided. Note that no examples were judged to fall into Category D.
Appendix I Evaluate Reasoning Fidelity
We evaluate the fidelity of reasoning in Section 5.3 and define a taxonomy for a deep analysis of reasoning generated by our guardrail.
Figure 17: Instruction for LLM/human annotation to evaluate fidelity of reasoning.